### v2

# VAEmo: Efficient Representation Learning for Visual-Audio Emotion with Knowledge Injection

Abstract.
Audiovisual emotion recognition (AVER) aims to infer human emotions from nonverbal visual-audio (VA) cues, offering modality-complementary and language-agnostic advantages. However, AVER remains challenging due to the inherent ambiguity of emotional expressions, cross-modal expressive disparities, and the scarcity of reliably annotated data. Recent self-supervised AVER approaches have introduced strong multimodal representations, yet they predominantly rely on modality-specific encoders and coarse content-level alignment, limiting fine-grained emotional semantic modeling. To address these issues, we propose VAEmo, an efficient two-stage framework for emotion-centric joint VA representation learning with external knowledge injection. In Stage 1, a unified and lightweight representation network is pre-trained on large-scale speaker-centric VA corpora via masked reconstruction and contrastive objectives, mitigating the modality gap and learning expressive, complementary representations without emotion labels. In Stage 2, multimodal large language models automatically generate detailed affective descriptions according to our well-designed chain-of-thought prompting for only a small subset of VA samples; these rich textual semantics are then injected by aligning their corresponding embeddings with VA representations through dual-path contrastive learning, further bridging the emotion gap. Extensive experiments on multiple downstream AVER benchmarks show that VAEmo achieves state-of-the-art performance with a compact design, highlighting the benefit of unified cross-modal encoding and emotion-aware semantic guidance for efficient, generalizable VA emotion representations. Source code and pre-trained models will be available at https://github.com/MSA-LMC/VAEmo.
1. Introduction
Understanding nuanced human emotion is a critical capability for next-generation AI systems that pursue natural and empathetic interactions with users (Liu et al., 2024; Bravo et al., 2025; Huang et al., 2019; Lazic et al., 2025; Magni et al., 2024). Human affect is inherently multimodal, expressed through vocal intonation, facial expressions, body postures, and linguistic content—each modality providing complementary cues for emotion analysis. This study focuses on the task of audio-visual emotion recognition (AVER), which aims to infer human affective states from audio-visual signals in a language-agnostic and modality-complementary manner (Mehrabian, 2017; Wu et al., 2014).
Emotion recognition is a fine-grained and semantically complex task, where the gap between general content features and emotion-specific features is often more challenging to bridge than the gap between modalities themselves. For instance, a person’s talking face might show sadness in one clip and anger in another, despite consistent identity and modality inputs, the emotional semantics diverge significantly. This indicates the necessity of learning more nuanced and comprehensive emotional representations. Unfortunately, emotional information is subtle, subjective, and difficult to precisely quantify and annotate at scale (Kollias and Zafeiriou, 2018; Li and Deng, 2022; Wang et al., 2022), hindering conventional supervised approaches and making it challenging to develop powerful multimodal emotion representations (Liu et al., 2022; Zhang et al., 2023b; Li et al., 2023; Chen et al., 2024b).
To overcome these limitations in emotion recognition, this field has increasingly turned to self-supervised learning (SSL) techniques (He et al., 2022; Tong et al., 2022; Huang et al., 2022; Radford et al., 2021; Hsu et al., 2021). By leveraging large amounts of unlabeled audio-visual data, SSL methods have shown great potential for learning rich multimodal representations without requiring expensive emotion annotations (Gong et al., [n. d.]; Huang et al., 2023; Zellers et al., 2021). However, existing SSL methods, while proficient at capturing general audio-visual correspondences, often overlook subtle emotional cues that require specialized attention. This creates a fundamental mismatch for emotion recognition, as these methods fail to target the extraction of emotion-specific features and fine-grained emotional semantic alignment during pre-training. Consequently, models derived from these self-supervised manner often miss emotional nuances, unable to effectively capture intra-modal emotional cues or align fine-grained emotional semantics across modalities in AVER downstream tasks.
Meanwhile, the emergence of Multimodal Large Language Models (MLLMs) (Xu et al., 2024) offers promising new directions for addressing these challenges. Although general large models still show moderate performance on representative emotion recognition tasks (Lian et al., 2024; Xing et al., 2024), they encode rich world knowledge and latent commonsense priors that are often neglected by common self-supervised approaches. Incorporating such knowledge during pre-training could significantly benefit emotion representation learning. Prior work like FineCLIPER (Chen et al., 2024a) and Emotion-LLaMA (Cheng et al., 2024) has begun exploring large models by generating captions to enhance model learning. However, FineCLIPER focuses solely on the visual modality, ignoring multimodal information, and operates during fine-tuning stage rather than pre-training stage. Its inference requires captions for additional supervision, limiting rapid transfer to downstream tasks. Similarly, Emotion-LLaMA only extracts peak frames from videos for facial action unit analysis, making its performance heavily reliant on the accuracy of peak frame selection, which is often limited in in-the-wild environments.
In response to these challenges, we introduce VAEmo, a novel framework specifically designed for audio-visual emotion representation learning with a two-stage training paradigm. Inspired by recent works on unified encoder designs for multimodal learning (Srivastava and Sharma, 2024; Bachmann et al., 2022; Zhang et al., 2023a), our VAEmo adopts a an efficient unified encoder that processes both audio and visual inputs, significantly reducing model parameters while maintaining excellent recognition performance . Unlike previous approaches that typically use separate encoders for each modality (dual-tower models), our unified encoder enables tighter integration of multimodal features, facilitating more efficient and effective representation learning. As illustrated in Fig. 1, our approach first establishes foundational cross-modal correspondences through contrastive and masked reconstruction learning (Stage 1) and then injects emotion-specific knowledge using MLLMs (Stage 2).
Our knowledge injection pipeline leverages specialized MLLMs to generate emotion-focused descriptions from both modalities, creating a rich semantic bridge between raw signals and emotional concepts. Through a dual-path contrastive learning approach that maintains separate audio-text and video-text learning paths while optimizing the same representation network, we preserve modality-specific characteristics while enriching emotional semantics. Our carefully designed two-stage training paradigm effectively addresses the gap between general audio-visual representation learning and emotion-specific feature extraction. By first establishing strong cross-modal representations in Stage 1 before injecting emotion-specific knowledge in Stage 2, we create a more stable learning trajectory. Extensive experiments demonstrate that VAEmo consistently outperforms previous state-of-the-art approaches while using significantly fewer parameters. In summary, our work makes the following contributions:
A novel two-stage training paradigm that progressively transitions from generic audio-visual representations into emotion-centric representations through contrastive learning, masked prediction, and knowledge injection.
An innovative knowledge injection method leveraging MLLMs to generate emotion-focused descriptions, creating a rich semantic bridge between raw signals and emotional concepts through a dual-path contrastive learning approach.
Comprehensive experimental validation revealing that for emotion-specific tasks, architectural simplicity coupled with targeted knowledge injection yields superior results compared to more complex models.
2. Related Work
Audio-Visual Emotion Recognition. Audio-visual emotion recognition (AVER) integrates speech and facial expression analysis, providing complementary multi-modal perspectives on emotion understanding (Mehrabian, 2017; Wu et al., 2014). Early approaches relied on supervised learning with hand-crafted features, such as eGeMAPS (Eyben et al., 2015) for audio and VGGFace (Cao et al., 2018) for visual features. With the advancement of deep learning, end-to-end approaches became prevalent. Liu et al. (Liu et al., 2022) proposed T-ESFL, which employs collaborative multi-task learning to improve AVER performance. Zhang et al. (Zhang et al., 2023b) developed T-MEP, introducing attention mechanisms to enhance cross-modal interaction. Recently, self-supervised approaches have made significant breakthroughs, such as AW-HuBERT (Tran et al., 2023) leveraged audio-visual self-supervised pre-training to enhance scene adaptation capabilities, and HiCMAE (Sun et al., 2024a) learned expressive representations through hierarchical contrastive masked autoencoding. Additionally, large models have been applied to emotion analysis. Chen et al. (Chen et al., 2024a) introduced FineCLIPER, which utilizes LLMs to generate emotional annotations, and Zhao et al. (Zhao and Patras, 2023) proposed DFER-CLIP, which explores visual prompt tuning for emotion recognition. In contrast to these methods, our VAEmo adopts a compact single-stream architecture and two-stage training paradigm, significantly enhancing emotional representation capacity while maintaining high efficiency.
Self-Supervised Learning and Knowledge Injection. Self-supervised learning (SSL) has become the dominant paradigm for representation learning. In the visual domain, MAE (He et al., 2022) employs high-ratio masked reconstruction pre-training, while Tong et al. (Tong et al., 2022) extended this to the spatiotemporal domain. For audio, Hsu et al.’s (Hsu et al., 2021) HuBERT and Chen et al.’s (Chen et al., 2022) WavLM utilize hidden unit prediction to establish powerful speech representations. In multi-modal learning, Gong et al. (Gong et al., [n. d.]) combined contrastive learning with MAE, while MAViL (Huang et al., 2023) explored more on contrastive object. Despite these advances, existing SSL approaches primarily focus on general-purpose representation learning, often overlooking the injection of domain-specific knowledge such as emotion semantics. For knowledge injection, with the emergence of large models, researchers have begun exploring LLMs as external knowledge sources. LLM2CLIP (Wu et al., [n. d.]) injects knowledge from large models into CLIP (Radford et al., 2021) models to enhance the performance of visual-language models. MATE (Jang et al., 2024) uses LoRA-tuned (Shen et al., [n. d.]) large language models to enhance the performance of CLIP visual encoders. Different from these works, our method uniquely uses multimodal LLMs (Qwen2-VL (Wang et al., 2024) and Qwen2-Audio (Chu et al., 2024)) to generate emotion-focused descriptions, incorporating emotional knowledge through dual-path contrastive learning that effectively bridges semantic gap between general and emotion-specific representations.
3. Approach
Problem Formulation. We focus on learning emotion representations from the vast amount of audio and video data available on the Internet. With audio and video modalities indexed by , each modality provides complementary emotional cues through different channels—speech patterns (e.g., tone, prosody) for audio and facial dynamics (e.g., expressions, micro-movements) for video. We propose a simple yet powerful model architecture coupled with a two-stage training framework to capture these emotional representations effectively.
Overview. As shown in Fig. 2, our method follows a two-stage training paradigm. Raw inputs are encoded through our representation network, which consists of three key components: modality-specific tokenizers, a feature transformation encoder, and a cross-modal fusion encoder. Based on different training stages, the encoded audio and visual features are input to stage-specific task heads. In Stage 1, this includes a combination of masked reconstruction and contrastive learning objectives, facilitating the establishment of cross-modal representations. In Stage 2, this becomes our knowledge injection framework, where captions generated by our carefully designed Chain of Thought (CoT)-based caption generation mechanism serve as emotional knowledge injected into representation network. The following sections provide detailed descriptions of VAEmo’s architecture and two-stage training paradigm.
3.1. Network Architecture
VAEmo architecture consists of four main components as shown in Fig. 2: (1) Modality-specific tokenizers : convert raw audio and visual inputs into token sequences with consistent dimensions; (2) A shared-parameter feature transformation encoder : encode each modality with multiple Transformer layers using shared parameters. (3) A cross-modal fusion encoder : integrates information by concatenating outputs from across modalities, followed by Transformer layers with shared parameters; (4) Stage-specific heads : vary depending on the training stage .
Both the feature transformation encoder and fusion encoder are implemented using a Vision Transformer (ViT)-based architecture. This intentionally simplified design offers two significant advantages: (1) it allows us to focus research attention on our novel two-stage training methodology rather than architectural innovations, providing clearer insights into the effectiveness of our knowledge injection approach; and (2) the unified architecture with shared parameters across modalities enables tighter integration of multimodal features in our encoder, while reducing the model’s parameter count compared to traditional dual-stream methods commonly used in audio-visual representation learning.
Data Processing: Given unlabeled video data with accompanying audio tracks , we demultiplex and pre-process the data into a video frame sequence and an audio log-mel spectrogram for the respective modalities. These raw inputs are transformed into token sequences and using modality-specific tokenizers as shown in Fig. 2. For audio, we treat mel-spectrograms as single-channel images and apply non-overlapping 2D convolutions (patch size ) to tokenization. For video, we employ 3D convolutions to extract spatiotemporal ”tubes” from frame sequences. This yields audio tokens and video tokens. This tokenization process ensures consistent output dimensions across modalities. The tokenizers thus produce dimension-aligned sequences , where is the dimension of the embedding.
Our model architecture consists of a unified representation network and stage-specific task heads .
where and represent the modality-specific tokenizers, represents the representation network parameters, and denotes token-wise concatenation of feature sequences. The data flow for a specific training stage can be expressed as:
The overall learning process involves optimizing both representation parameters and Stage-specific parameters :
where and denotes the loss function for training stage . While the task heads change across different stages, they all optimize the same underlying representation network , allowing us to effectively transfer knowledge across stages.
Stage-Specific Task Heads: Our framework employs three distinct task heads across training stages:
Stage 1: A dual-purpose head () supporting both masked reconstruction and contrastive learning objectives. The reconstruction component decodes masked tokens while the contrastive component aligns audio-visual representation.
Stage 2: A knowledge injection head () that facilitates dual-path contrastive learning between audio-text and video-text pairs using MLLM-generated descriptions.
Fine-tuning: A simple classification head () implemented as a single-layer MLP that projects the fused representations to emotion categories or dimensional scores.
3.2. Training Paradigm
Stage 1: Contrastive Learning and Masked Reconstruction. For Stage 1 pre-training, we employ a dual strategy combining contrastive learning and masked self-supervised learning, an approach proven effective in prior works like CAV-MAE (Gong et al., [n. d.]) and MAViL (Huang et al., 2023). Our contrastive objective employs the standard InfoNCE loss with temperature scaling to align the global representations of paired video and audio. For each video–audio pair, the global representations for video and audio are obtained by token-wise mean pooling over the output feature sequences and , respectively. Simultaneously, we perform masked self-supervised learning by randomly masking tokens (80% for audio, 90% for video) and tasking the model to reconstruct the original data from these partially masked inputs. The reconstruction process involves processing masked inputs through our encoders, fusing modality-specific features, and computing the mean squared error loss between original and reconstructed data. This combined approach offers complementary benefits: contrastive learning captures high-level cross-modal correspondences, while reconstruction preserves fine-grained details within each modality. After Stage 1 pre-training, we discard the reconstruction decoder and retain only the representation network for Stage 2, allowing the model to first learn generalizable speech-face representations before incorporating emotional knowledge.
Stage 2: Knowledge Injection with MLLMs. In the second stage, we further enhance our model’s emotion representation capabilities based on the stage 1 parameters through knowledge injection. We develop a fully automated knowledge extraction and injection pipeline leveraging Multimodal Large Language Models (MLLMs). Using modality-specific MLLMs—Qwen2-VL for video and Qwen2-Audio for audio, we generate rich emotional descriptions through a carefully designed Chain-of-Thought (CoT) prompting strategy.
Our CoT process first guides MLLMs to describe observable physical characteristics—facial action units, overall posture, and body movements for video, or acoustic properties for audio. Based on these objective features, MLLMs then assess emotional states, including intensity (arousal) and valence, and finally analyze overall emotional changes throughout the sequence. To ensure that the generated captions are both factual and emotionally relevant, we introduce a refinement mechanism. First, we explicitly instruct the MLLMs to ground all inferences in observable features. Then, we conduct multiple inference passes per input and apply a majority voting strategy to filter out inconsistent or hallucinated outputs. We employ a secondary LLM (GPT-4o-mini (Hurst et al., 2024)) to filter the results by discarding emotion-irrelevant content.This refinement pipeline encourages the generation of expressive and specialized affective vocabulary (e.g., “assertiveness,” “amusement,” “demeanor”), enabling the model to capture subtle and compound emotional states. As illustrated in Fig. 3, these high-quality descriptions serve as valuable supervision signals during training.
Since the generated captions are much longer than CLIP (Radford et al., 2021)’s 77-token limit (in our implementation, tokens counted using GPT-2 (Radford et al., [n. d.]) tokenizer can reach up to 100) and contain specialized emotional vocabulary absent from CLIP’s training data, we employ LLM-based embeddings for text representation. Specifically, we use the NV-Embed-v2 (Lee et al., 2024) model, which adapts a decoder-only LLM as an embedding generator by employing specialized latent attention layers and removing causal attention masks. Our ablation experiments demonstrate that LLM embeddings significantly outperform both CLIP text embeddings and BERT embeddings for our emotion recognition tasks. After encoding the captions into text embeddings, we further process each embedding through a modality-specific adapter module, resulting in and for audio and video, respectively. The adapter design follows the implementation in LLM2CLIP (Wu et al., [n. d.]). We then implement a dual-path contrastive learning approach (Fig. 2), which replaces the traditional audio-visual contrastive objective from Stage 1. Specifically, we construct two parallel contrastive objectives: one between audio caption embedding and audio global representation, and the other between video caption embedding and video global representation.
where represents either audio or video. The total loss for Stage 2 is a weighted combination of these components:
A crucial aspect of this dual-path approach is that both objectives simultaneously optimize the same encoder through different gradient paths. and are the token-wise mean-pooled embedding of the feature transformation encoder output for audio and video, respectively. The gradients from the two contrastive losses flow back to update the same set of parameters. This creates a rich training signal that enhances the encoder’s ability to aware emotional knowledge for both modalities.
After completing both training stages, we extract the encoder and fine-tune it on downstream tasks to leverage the emotion-rich representations for specific applications. The sequential nature of our training paradigm—from general audio-visual correspondences to emotion-specific knowledge—creates a robust foundation for visual-audio emotion recognition.
4. Experiments
In this section, we evaluate the effectiveness of VAEmo on a suite of audio-visual emotion recognition (AVER) tasks. Our experiments cover both categorical and dimensional emotion recognition paradigms across diverse datasets, spanning from controlled lab environments to unconstrained in-the-wild scenarios.
4.1. Experimental Setup
| Dataset | Setting | Task | #Classes | Samples | Metric |
| VoxCeleb2 (Chung et al., 2018) | - | Pre-training | - | 1,360,531 | - |
| MAFW (Liu et al., 2022) | Wild | Categorical | 11 | 9,172 | UAR, WAR |
| DFEW (Jiang et al., 2020) | Wild | Categorical | 7 | 11,697 | UAR, WAR |
| CREMA-D (Cao et al., 2014) | Lab | Categorical | 6 | 7,442 | UAR, WAR |
| MSP-IMPROV (Busso et al., 2016) | Lab | Categorical | 4 | 7,798 | UAR, WAR |
| Werewolf-XL (Zhang et al., 2021) | Lab | Dimensional | 3 | 8,542 | PCC |
Datasets. Tab. 1 provides a comprehensive overview of all datasets used in our experiments; all used datasets are video data with accompanying audio tracks. For pre-training (including both Stage 1 and Stage 2), we use VoxCeleb2, which contains over 1 million utterances from 6,112 celebrities extracted from YouTube videos. This dataset provides a rich source of speech-face pairs for learning audio-visual representations. For downstream evaluation, we select representative datasets covering different aspects of emotion recognition: MAFW and DFEW for in-the-wild categorical emotion recognition, which present challenging scenarios with natural variations in lighting, pose, and acoustic conditions; CREMA-D and MSP-IMPROV for lab-controlled categorical emotion recognition with high-quality recordings under controlled settings; and Werewolf-XL for dimensional emotion recognition, which annotates emotions along continuous dimensions including valence, arousal, and dominance.
Evaluation Metrics. For categorical emotion recognition, we report both Unweighted Average Recall (UAR) and Weighted Average Recall (WAR) to address potential class imbalance issues. For dimensional emotion recognition, we use the Pearson Correlation Coefficient (PCC) following standard practice. For those datasets using cross-validation, we combine the predictions and labels from all folds to calculate the final UAR and WAR, the same as the evaluation in (Sun et al., 2024a).
4.2. Implementation Details
Preprocessing. For video preprocessing, we employ different strategies during pre-training and fine-tuning. During pre-training, we follow (Sun et al., 2023) to crop each frame to pixels and sample 16 frames to form video clips with dimensions (frames height width channels), which is computationally efficient for large-scale processing. For fine-tuning on downstream tasks, we employ InsightFace (DeepInsight, 2018) to detect and crop facial regions more precisely from each frame. The facial crops are resized to pixels. For audio, we extract 128-dimensional mel-spectrograms with a 25 ms window size and 10 ms shift, which are then converted to single-channel ”images” for processing, following (Sun et al., 2024a).
Model Configuration. Our VAEmo framework employs a ViT-Small architecture. The model maintains a consistent embedding dimension of 512 across all components with 8 attention heads. Our streamlined architecture consists of a feature transformation encoder which is 10-layer transformer followed by a fusion encoder which is 2-layer transformer, resulting in a total parameter count of 39M—significantly smaller than comparable models like HiCMAE-B (81M) and AV-HuBERT variants (103M).
Training Details. Our training proceeds in two distinct stages. For Stage 1 Pre-training, we train the model on the full VoxCeleb2 dataset using 8 NVIDIA A800 GPUs with a batch size of 170 per GPU. We use the AdamW optimizer with a base learning rate of 1.2e-3 and a cosine learning rate schedule. This stage involves 100 epochs of training, requiring approximately 48 hours to complete. For Stage 2 Knowledge Injection, to incorporate emotional knowledge, we randomly sample 10% of the VoxCeleb2 dataset (approximately 100K samples) and generate emotion-focused captions using our MLLM pipeline. To prevent catastrophic forgetting, we employ LayerNorm tuning (Zhao et al., [n. d.]) where only LayerNorm parameters are updated while keeping other weights frozen. This stage involves 10 epochs of training on the captioned subset.
4.3. Comparison with State-of-the-Art Methods
| Method | SSL | Modality | #Params(M) | UAR | WAR |
| Audio-only approaches: | |||||
| HuBERT (Hsu et al., 2021) | ✓ | A | 95 | 25.00 | 32.60 |
| WavLM-Plus (Chen et al., 2022) | ✓ | A | 95 | 26.33 | 34.07 |
| Visual-only approaches: | |||||
| DFER-CLIP (Zhao and Patras, 2023) | ✓ | V | 153 | 39.89 | 52.55 |
| SVFAP (Sun et al., 2024b) | ✓ | V | 78 | 41.19 | 54.28 |
| MAE-DFER (Sun et al., 2023) | ✓ | V | 85 | 41.62 | 54.31 |
| S4D (Chen et al., 2025) | ✓ | V | 101 | 43.72 | 58.44 |
| Multi-modal approaches: | |||||
| T-ESFL (Liu et al., 2022) | A+V | - | 33.35 | 48.70 | |
| T-MEP (Zhang et al., 2023b) | A+V | 61 | 37.17 | 51.15 | |
| MMA-DFER (Chumachenko et al., 2024) | ✓ | A+V | - | 44.25 | 58.45 |
| HiCMAE-S (Sun et al., 2024a) | ✓ | A+V | 46 | 41.66 | 54.45 |
| HiCMAE-B (Sun et al., 2024a) | ✓ | A+V | 81 | 42.65 | 56.17 |
| VAEmo (ours) | ✓ | A+V | 39 | 45.67 | 58.91 |
| FineCLIPER (Chen et al., 2024a) | ✓ | T+V | N/A | 45.01 | 56.91 |
| Method | SSL | Mod. | #Params(M) | UAR | WAR |
| (A) Approaches using external pre-trained weights or additional labels: | |||||
| S2D (Chen et al., 2024b) | ✓ | V | N/A | 65.45 | 74.81 |
| S4D (Chen et al., 2025) | ✓ | V | 101 | 66.80 | 76.68 |
| FineCLIPER (Chen et al., 2024a) | ✓ | T+V | N/A | 65.98 | 76.21 |
| (B) Approaches trained from scratch on VoxCeleb2 only: | |||||
| MAE-DFER (Sun et al., 2023) | ✓ | V | 85 | 63.41 | 74.43 |
| HiCMAE-S (Sun et al., 2024a) | ✓ | A+V | 46 | 63.05 | 74.33 |
| HiCMAE-B (Sun et al., 2024a) | ✓ | A+V | 81 | 63.76 | 75.01 |
| VAEmo (ours) | ✓ | A+V | 39 | 64.02 | 75.78 |
| AMH (Yoon et al., 2020) | A+V | - | 54.48 | 66.51 | |
| Method | SSL | Mod. | #Params(M) | UAR | WAR |
| HuBERT (Hsu et al., 2021) | ✓ | A | 95 | 72.72 | 72.57 |
| WavLM-Plus (Chen et al., 2022) | ✓ | A | 95 | 73.34 | 73.39 |
| SVFAP (Sun et al., 2024b) | ✓ | V | 78 | 77.31 | 77.37 |
| MAE-DFER (Sun et al., 2023) | ✓ | V | 85 | 77.33 | 77.38 |
| VQ-MAE-AV (Hazmoune and Bougamouza, 2024) | ✓ | A+V | 30 | - | 80.40 |
| HiCMAE-S (Sun et al., 2024a) | ✓ | A+V | 46 | 84.46 | 84.38 |
| HiCMAE-B (Sun et al., 2024a) | ✓ | A+V | 81 | 84.91 | 84.89 |
| VAEmo (ours) | ✓ | A+V | 39 | 85.71 | 85.68 |
| Method | SSL | Mod. | #Params(M) | UAR | WAR |
| FAV-HuBERT (Tran et al., 2023) | ✓ | A+V | 103 | 61.05 | 68.35 |
| TAPT-HuBERT (Tran et al., 2023) | ✓ | A+V | 103 | 63.95 | 70.46 |
| AW-HuBERT (Tran et al., 2023) | ✓ | A+V | 103 | 65.72 | 71.80 |
| HiCMAE-S (Sun et al., 2024a) | ✓ | A+V | 46 | 63.90 | 74.35 |
| HiCMAE-B (Sun et al., 2024a) | ✓ | A+V | 81 | 65.78 | 74.95 |
| VAEmo (ours) | ✓ | A+V | 39 | 66.12 | 76.79 |
| Method | SSL | Mod. | #Params(M) | Arousal | Valence | Dominance |
| eGeMAPS (Eyben et al., 2015) | A | - | 23.45 | 8.08 | 31.15 | |
| VGGFace (Cao et al., 2018) | V | - | 7.24 | 62.96 | 14.30 | |
| SVFAP (Sun et al., 2024b) | ✓ | V | 78 | 23.51 | 67.11 | 34.61 |
| HiCMAE (Sun et al., 2024a) | ✓ | A+V | 81 | 33.74 | 69.23 | 40.66 |
| VAEmo (ours) | ✓ | A+V | 39 | 39.80 | 69.72 | 52.87 |
Tabs. 2-6 present comprehensive results comparing our VAEmo framework with state-of-the-art approaches across five diverse emotion recognition benchmarks. Our model demonstrates superior performance while maintaining significantly fewer parameters (39M) than previous approaches (81 M- 103 M). Note that for approaches marked with N/A in parameter count (S2D and FineCLIPER), only fine-tunable parameters (9M and 20M, respectively) were reported in the original papers without specifying the total parameter count. Based on their use of ViT-base as the backbone, we estimate their total parameter count to exceed 86 M.
As shown in Tab. 2, on the challenging in-the-wild categorical emotion recognition dataset MAFW, our approach achieves 45.67% UAR and 58.91% WAR, outperforming the previous best method (MMA-DFER (Chumachenko et al., 2024)), which achieved 44.25% UAR and 58.45% WAR. On the DFEW dataset, while S4D (Chen et al., 2025) achieves higher UAR (66.80%), our method reaches 75.78% WAR, exceeding HiCMAE-B (Sun et al., 2024a) (75.01% WAR) with only 39M parameters compared to HiCMAE-B’s 81M and S4D’s 101M parameters. It is worth noting that S4D leverages additional static facial expression datasets for pre-training, while FineCLIPER incorporates extra modalities, including facial landmarks, and S2D similarly utilizes external data resources. Meanwhile, among all approaches that use only VoxCeleb2 for pre-training (as shown in Tab. 3, group B), VAEmo consistently achieves the best performance while requiring significantly fewer parameters, highlighting the effectiveness of our two-stage training paradigm and knowledge injection approach.
On the lab-controlled categorical emotion recognition dataset CREMA-D, our method achieves 85.71% UAR and 85.68% WAR, setting a new benchmark that surpasses larger models like HiCMAE-B (Sun et al., 2024a) (84.91% UAR, 84.89% WAR). On MSP-IMPROV, we obtain 66.12% UAR and 76.79% WAR, outperforming the previous best approach AW-HuBERT (Tran et al., 2023) (65.72% UAR, 71.80% WAR) by +0.40% UAR and +4.99% WAR while using 62% fewer parameters. These results demonstrate our model’s exceptional capability in controlled recording environments. Our method also excels at continuous emotion recognition on the Werewolf-XL dataset, achieving 39.80%, 69.72%, and 52.87% PCC for arousal, valence, and dominance dimensions, respectively. The improvements compared to HiCMAE (Sun et al., 2024a) are particularly substantial for arousal (+6.06%) and dominance (+12.21%), indicating our approach’s effectiveness at capturing subtle emotional intensity variations. These results collectively demonstrate that architectural simplicity coupled with our two-stage knowledge injection paradigm creates a powerful foundation for emotion recognition across diverse settings.
Furthermore, as shown in Fig. 4, our proposed model, VAEmo (39M parameters), outperforms other state-of-the-art approaches on the MAFW dataset with a significantly lower parameter count. Specifically, VAEmo achieves a WAR of 58.91%, while models such as HiCMAE-B (81M parameters) and S4D (101M parameters) perform at 56.17% and 58.44%, respectively.
4.4. Ablation Studies
A bar chart comparing the performance metrics (UAR and WAR) between Stage 1 and Stage 2 of the VAEmo framework on the first fold of the MAFW dataset, showing consistent improvements after knowledge injection.
Effectiveness of the Proposed Two-stage Training.
| Training Strategy | MAFW | MSP-IMPROV | Werewolf-XL | ||||
| UAR | WAR | UAR | WAR | A | V | D | |
| Stage 1 only | 43.07 | 57.69 | 65.53 | 75.67 | 35.51 | 69.48 | 43.91 |
| Two-stage | 45.67 | 58.91 | 66.12 | 76.79 | 39.80 | 69.72 | 52.87 |
| Improvement | +2.60 | +1.22 | +0.59 | +1.12 | +4.29 | +0.24 | +8.96 |
| Joint training | 41.25 | 56.43 | 64.87 | 74.92 | 33.82 | 68.95 | 41.73 |
First, we investigate the effectiveness of our two-stage training approach by comparing three training strategies: (1) Stage 1 only (contrastive learning and masked self-supervised learning), (2) our full two-stage approach (Stage 1 followed by Stage 2), and (3) joint training (applying both objectives simultaneously). As shown in Tab. 7, the sequential two-stage approach consistently outperforms both alternatives. Joint training underperforms Stage 1-only training, suggesting that introducing MLLM knowledge without establishing strong cross-modal representations may confuse the learning process. This confirms our hypothesis that establishing general audio-visual correspondences before injecting emotion-specific knowledge creates a more effective foundation for emotion recognition. A significant finding is that even our Stage 1 model (without knowledge injection) already outperforms most existing approaches across datasets, as shown in Tab. 7. The additional knowledge injection in Stage 2 provides consistent improvements, with gains of +2.60% UAR on MAFW (from 43.07% to 45.67%) and +0.59% UAR on MSP-IMPROV (from 65.53% to 66.12%). This confirms that our architecture with targeted MLLM knowledge integration effectively captures nuanced emotional cues that more complex models might miss. Furthermore, the t-SNE visualization in Fig. 5 clearly shows how the knowledge injection in Stage 2 produces more distinct emotion clusters than the representations after Stage 1.
Analysis of Knowledge Injection Strategies.
| Knowledge Injection Strategy | UAR | WAR |
| Embedding Type: | ||
| CLIP embedding | 43.21 | 58.20 |
| BERT embedding | 44.35 | 58.14 |
| LLM embedding | 45.67 | 58.91 |
| Caption Generation: | ||
| Basic MLLM prompt | 43.41 | 57.63 |
| Emotion-focused prompt | 43.32 | 58.45 |
| Chain-of-Thought prompt | 45.67 | 58.91 |
As described in Section 3.2, we use CoT prompting for caption generation and LLM for caption embedding. Next, we examine various aspects of our knowledge injection approach in Stage 2. Tab. 8 shows the impact of different embedding types, caption generation approaches, and contrastive learning objectives. For embedding types, LLM-based embeddings outperform both CLIP embeddings and BERT embeddings, primarily because our captions exceed CLIP’s 77-token length limit and contain specialized emotional vocabulary requiring the richer semantic understanding of LLMs. For caption generation, Chain-of-Thought prompting represents one of our key innovations, producing significantly richer emotion-relevant semantics than basic prompting approaches, while emotion-focused prompts offer moderate improvements over basic prompts, they lack the comprehensive emotional reasoning that CoT provides by first analyzing observable features before inferring emotional states.
Model Architecture Choices. To better clarify our architectural design decisions, we compare single-stream versus dual-stream approaches. For fair comparison, all experiments were conducted using 100 epochs of Stage 1 training. As shown in Tab. 9, our single-stream architecture achieves comparable performance while using 45% fewer parameters (39M vs. 71M). The dual-stream model essentially represents a simplified version of HiCMAE without hierarchical connections and cross-attention interactions. The single-stream model’s efficiency comes from sharing parameters across modalities, creating a unified parameter space for audio and visual processing. The shared space enables tighter integration where features are processed through identical transformations. This reveals that for emotion recognition, parameter sharing from early processing stages can lead to better generalization, even with simpler fusion methods. Although computation costs are similar, the smaller parameter size makes single-stream model easier to deploy.
| Architecture | UAR | WAR | #Params (M) | FLOPs (G) |
| Dual-stream | 43.01 | 56.67 | 71 | 49.33 |
| Single-stream | 43.07 | 57.69 | 39 | 49.33 |
| A-T vs. V-T weight: | UAR | WAR |
| 44.32 | 57.85 | |
| 45.35 | 58.64 | |
| 45.20 | 58.42 | |
| 45.67 | 58.91 | |
| 44.95 | 58.42 |
Hyper-parameter Analysis. Tab. 10 shows the effect of different audio-text vs. video-text weights in Stage 2. Slightly favoring audio () yields optimal performance, suggesting audio provides stronger emotional cues on MAFW. Extreme weighting in either direction reduces performance, confirming the importance of balanced modality integration during knowledge injection.
5. Conclusion
In this paper, we presented VAEmo, an efficient framework for audio-visual emotion representation learning that addresses the critical mismatch between general and emotion-specific representations. Through a carefully orchestrated two-stage pre-training approach on unlabeled audio-visual data, combined with our carefully designed emotional knowledge injection paradigm, our model effectively bridges the semantic gap between general audio-visual perception and emotion-specific understanding. Extensive experiments across multiple downstream benchmarks demonstrate that the VAEmo achieves state-of-the-art performance with strong generalization and a compact architecture.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (NSFC, No. 62202139, No. 62172138) and the Fundamental Research Funds for the Central Universities (No. JZ2025HGTB0226, No. JZ2024HGTG0310). The computation is completed on the HPC Platform of Hefei University of Technology.References
- (1)
- Bachmann et al. (2022) Roman Bachmann, David Mizrahi, Andrei Atanov, and Amir Zamir. 2022. Multimae: Multi-modal multi-task masked autoencoders. In European Conference on Computer Vision. Springer, 348–367.
- Bravo et al. (2025) Luis Bravo, Ciro Rodriguez, Pedro Hidalgo, and Cesar Angulo. 2025. A Systematic Review on Artificial Intelligence-Based Multimodal Dialogue Systems Capable of Emotion Recognition. Multimodal Technologies and Interaction 9, 3 (2025), 28.
- Busso et al. (2016) Carlos Busso, Srinivas Parthasarathy, Alec Burmania, Mohammed AbdelWahab, Najmeh Sadoughi, and Emily Mower Provost. 2016. MSP-IMPROV: An acted corpus of dyadic interactions to study emotion perception. IEEE Transactions on Affective Computing 8, 1 (2016), 67–80.
- Cao et al. (2014) Houwei Cao, David G Cooper, Michael K Keutmann, Ruben C Gur, Ani Nenkova, and Ragini Verma. 2014. Crema-d: Crowd-sourced emotional multimodal actors dataset. IEEE transactions on affective computing 5, 4 (2014), 377–390.
- Cao et al. (2018) Qiong Cao, Li Shen, Weidi Xie, Omkar M Parkhi, and Andrew Zisserman. 2018. Vggface2: A dataset for recognising faces across pose and age. In 2018 13th IEEE international conference on automatic face & gesture recognition (FG 2018). IEEE, 67–74.
- Chen et al. (2024a) Haodong Chen, Haojian Huang, Junhao Dong, Mingzhe Zheng, and Dian Shao. 2024a. Finecliper: Multi-modal fine-grained clip for dynamic facial expression recognition with adapters. In Proceedings of the 32nd ACM International Conference on Multimedia. 2301–2310.
- Chen et al. (2022) Sanyuan Chen, Chengyi Wang, Zhengyang Chen, Yu Wu, Shujie Liu, Zhuo Chen, Jinyu Li, Naoyuki Kanda, Takuya Yoshioka, Xiong Xiao, et al. 2022. Wavlm: Large-scale self-supervised pre-training for full stack speech processing. IEEE Journal of Selected Topics in Signal Processing 16, 6 (2022), 1505–1518.
- Chen et al. (2024b) Yin Chen, Jia Li, Shiguang Shan, Meng Wang, and Richang Hong. 2024b. From static to dynamic: Adapting landmark-aware image models for facial expression recognition in videos. IEEE Transactions on Affective Computing (2024).
- Chen et al. (2025) Yin Chen, Jia Li, Yu Zhang, Zhenzhen Hu, Shiguang Shan, Meng Wang, and Richang Hong. 2025. Static for Dynamic: Towards a Deeper Understanding of Dynamic Facial Expressions Using Static Expression Data. arXiv:2409.06154 [cs.CV] https://arxiv.org/abs/2409.06154
- Cheng et al. (2024) Zebang Cheng, Zhi-Qi Cheng, Jun-Yan He, Kai Wang, Yuxiang Lin, Zheng Lian, Xiaojiang Peng, and Alexander Hauptmann. 2024. Emotion-llama: Multimodal emotion recognition and reasoning with instruction tuning. Advances in Neural Information Processing Systems 37 (2024), 110805–110853.
- Chu et al. (2024) Yunfei Chu, Jin Xu, Qian Yang, Haojie Wei, Xipin Wei, Zhifang Guo, Yichong Leng, Yuanjun Lv, Jinzheng He, Junyang Lin, et al. 2024. Qwen2-audio technical report. arXiv preprint arXiv:2407.10759 (2024).
- Chumachenko et al. (2024) Kateryna Chumachenko, Alexandros Iosifidis, and Moncef Gabbouj. 2024. MMA-DFER: MultiModal Adaptation of unimodal models for Dynamic Facial Expression Recognition in-the-wild. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 4673–4682.
- Chung et al. (2018) Joon Son Chung, Arsha Nagrani, and Andrew Zisserman. 2018. Voxceleb2: Deep speaker recognition. arXiv preprint arXiv:1806.05622 (2018).
- DeepInsight (2018) DeepInsight. 2018. InsightFace: 2D and 3D Face Analysis Project. https://github.com/deepinsight/insightface.
- Eyben et al. (2015) Florian Eyben, Klaus R Scherer, Björn W Schuller, Johan Sundberg, Elisabeth André, Carlos Busso, Laurence Y Devillers, Julien Epps, Petri Laukka, Shrikanth S Narayanan, et al. 2015. The Geneva minimalistic acoustic parameter set (GeMAPS) for voice research and affective computing. IEEE transactions on affective computing 7, 2 (2015), 190–202.
- Gong et al. ([n. d.]) Yuan Gong, Andrew Rouditchenko, Alexander H Liu, David Harwath, Leonid Karlinsky, Hilde Kuehne, and James R Glass. [n. d.]. Contrastive Audio-Visual Masked Autoencoder. In The Eleventh International Conference on Learning Representations.
- Hazmoune and Bougamouza (2024) Samira Hazmoune and Fateh Bougamouza. 2024. Using transformers for multimodal emotion recognition: Taxonomies and state of the art review. Engineering Applications of Artificial Intelligence 133 (2024), 108339.
- He et al. (2022) Kaiming He, Xinlei Chen, Saining Xie, Yanghao Li, Piotr Dollár, and Ross Girshick. 2022. Masked autoencoders are scalable vision learners. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 16000–16009.
- Hsu et al. (2021) Wei-Ning Hsu, Benjamin Bolte, Yao-Hung Hubert Tsai, Kushal Lakhotia, Ruslan Salakhutdinov, and Abdelrahman Mohamed. 2021. Hubert: Self-supervised speech representation learning by masked prediction of hidden units. IEEE/ACM transactions on audio, speech, and language processing 29 (2021), 3451–3460.
- Huang et al. (2019) Ming-Hui Huang, Roland Rust, and Vojislav Maksimovic. 2019. The feeling economy: Managing in the next generation of artificial intelligence (AI). California management review 61, 4 (2019), 43–65.
- Huang et al. (2023) Po-Yao Huang, Vasu Sharma, Hu Xu, Chaitanya Ryali, Yanghao Li, Shang-Wen Li, Gargi Ghosh, Jitendra Malik, Christoph Feichtenhofer, et al. 2023. Mavil: Masked audio-video learners. Advances in Neural Information Processing Systems 36 (2023), 20371–20393.
- Huang et al. (2022) Po-Yao Huang, Hu Xu, Juncheng Li, Alexei Baevski, Michael Auli, Wojciech Galuba, Florian Metze, and Christoph Feichtenhofer. 2022. Masked autoencoders that listen. Advances in Neural Information Processing Systems 35 (2022), 28708–28720.
- Hurst et al. (2024) Aaron Hurst, Adam Lerer, Adam P Goucher, Adam Perelman, Aditya Ramesh, Aidan Clark, AJ Ostrow, Akila Welihinda, Alan Hayes, Alec Radford, et al. 2024. Gpt-4o system card. arXiv preprint arXiv:2410.21276 (2024).
- Jang et al. (2024) Young Kyun Jang, Junmo Kang, Yong Jae Lee, and Donghyun Kim. 2024. MATE: Meet At The Embedding–Connecting Images with Long Texts. arXiv preprint arXiv:2407.09541 (2024).
- Jiang et al. (2020) Xingxun Jiang, Yuan Zong, Wenming Zheng, Chuangao Tang, Wanchuang Xia, Cheng Lu, and Jiateng Liu. 2020. Dfew: A large-scale database for recognizing dynamic facial expressions in the wild. In Proceedings of the 28th ACM international conference on multimedia. 2881–2889.
- Kollias and Zafeiriou (2018) Dimitrios Kollias and Stefanos Zafeiriou. 2018. Aff-wild2: Extending the aff-wild database for affect recognition. arXiv preprint arXiv:1811.07770 (2018).
- Lazic et al. (2025) Milan Lazic, Earl Woodruff, and Jenny Jun. 2025. The Next Generation of Personalized Educational AI: Integrating Emotion and Cognition. In Social Robots in Education: How to Effectively Introduce Social Robots into Classrooms. Springer, 63–82.
- Lee et al. (2024) Chankyu Lee, Rajarshi Roy, Mengyao Xu, Jonathan Raiman, Mohammad Shoeybi, Bryan Catanzaro, and Wei Ping. 2024. Nv-embed: Improved techniques for training llms as generalist embedding models. arXiv preprint arXiv:2405.17428 (2024).
- Li et al. (2023) Hanting Li, Hongjing Niu, Zhaoqing Zhu, and Feng Zhao. 2023. Intensity-aware loss for dynamic facial expression recognition in the wild. In Proceedings of the AAAI conference on artificial intelligence, Vol. 37. 67–75.
- Li and Deng (2022) Shan Li and Weihong Deng. 2022. Deep Facial Expression Recognition: A Survey. IEEE Transactions on Affective Computing 13, 3 (2022), 1195–1215. doi:10.1109/TAFFC.2020.2981446
- Lian et al. (2024) Zheng Lian, Licai Sun, Haiyang Sun, Kang Chen, Zhuofan Wen, Hao Gu, Bin Liu, and Jianhua Tao. 2024. Gpt-4v with emotion: A zero-shot benchmark for generalized emotion recognition. Information Fusion 108 (2024), 102367.
- Liu et al. (2022) Yuanyuan Liu, Wei Dai, Chuanxu Feng, Wenbin Wang, Guanghao Yin, Jiabei Zeng, and Shiguang Shan. 2022. Mafw: A large-scale, multi-modal, compound affective database for dynamic facial expression recognition in the wild. In Proceedings of the 30th ACM international conference on multimedia. 24–32.
- Liu et al. (2024) Yang Liu, Janne Kauttonen, Bowen Zhao, Xiaobai Li, and Wei Peng. 2024. Towards Emotion AI to next generation healthcare and education. 1533053 pages.
- Magni et al. (2024) Domitilla Magni, Giovanna Del Gaudio, Armando Papa, and Valentina Della Corte. 2024. Digital humanism and artificial intelligence: the role of emotions beyond the human–machine interaction in Society 5.0. Journal of Management History 30, 2 (2024), 195–218.
- Mehrabian (2017) Albert Mehrabian. 2017. Communication without words. In Communication theory. Routledge, 193–200.
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In International conference on machine learning. PmLR, 8748–8763.
- Radford et al. ([n. d.]) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. [n. d.]. Language models are unsupervised multitask learners. ([n. d.]).
- Shen et al. ([n. d.]) Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, et al. [n. d.]. Lora: Low-rank adaptation of large language models. ([n. d.]).
- Srivastava and Sharma (2024) Siddharth Srivastava and Gaurav Sharma. 2024. Omnivec2-a novel transformer based network for large scale multimodal and multitask learning. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 27412–27424.
- Sun et al. (2023) Licai Sun, Zheng Lian, Bin Liu, and Jianhua Tao. 2023. Mae-dfer: Efficient masked autoencoder for self-supervised dynamic facial expression recognition. In Proceedings of the 31st ACM International Conference on Multimedia. 6110–6121.
- Sun et al. (2024a) Licai Sun, Zheng Lian, Bin Liu, and Jianhua Tao. 2024a. Hicmae: Hierarchical contrastive masked autoencoder for self-supervised audio-visual emotion recognition. Information Fusion 108 (2024), 102382.
- Sun et al. (2024b) Licai Sun, Zheng Lian, Kexin Wang, Yu He, Mingyu Xu, Haiyang Sun, Bin Liu, and Jianhua Tao. 2024b. SVFAP: Self-supervised video facial affect perceiver. IEEE Transactions on Affective Computing (2024).
- Tong et al. (2022) Zhan Tong, Yibing Song, Jue Wang, and Limin Wang. 2022. Videomae: Masked autoencoders are data-efficient learners for self-supervised video pre-training. Advances in neural information processing systems 35 (2022), 10078–10093.
- Tran et al. (2023) Minh Tran, Yelin Kim, Che-Chun Su, Cheng-Hao Kuo, and Mohammad Soleymani. 2023. Saaml: A framework for semi-supervised affective adaptation via metric learning. In Proceedings of the 31st ACM International Conference on Multimedia. 6004–6015.
- Wang et al. (2024) Peng Wang, Shuai Bai, Sinan Tan, Shijie Wang, Zhihao Fan, Jinze Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, et al. 2024. Qwen2-vl: Enhancing vision-language model’s perception of the world at any resolution. arXiv preprint arXiv:2409.12191 (2024).
- Wang et al. (2022) Yan Wang, Yixuan Sun, Yiwen Huang, Zhongying Liu, Shuyong Gao, Wei Zhang, Weifeng Ge, and Wenqiang Zhang. 2022. Ferv39k: A large-scale multi-scene dataset for facial expression recognition in videos. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 20922–20931.
- Wu et al. ([n. d.]) Aoqi Wu, Yifan Yang, Xufang Luo, Yuqing Yang, Chunyu Wang, Liang Hu, Xiyang Dai, Dongdong Chen, Chong Luo, Lili Qiu, et al. [n. d.]. LLM2CLIP: Powerful Language Model Unlock Richer Visual Representation. In NeurIPS 2024 Workshop: Self-Supervised Learning-Theory and Practice.
- Wu et al. (2014) Chung-Hsien Wu, Jen-Chun Lin, and Wen-Li Wei. 2014. Survey on audiovisual emotion recognition: databases, features, and data fusion strategies. APSIPA transactions on signal and information processing 3 (2014), e12.
- Xing et al. (2024) Bohao Xing, Zitong Yu, Xin Liu, Kaishen Yuan, Qilang Ye, Weicheng Xie, Huanjing Yue, Jingyu Yang, and Heikki Kälviäinen. 2024. Emo-llama: Enhancing facial emotion understanding with instruction tuning. arXiv preprint arXiv:2408.11424 (2024).
- Xu et al. (2024) Yaoxun Xu, Hangting Chen, Jianwei Yu, Qiaochu Huang, Zhiyong Wu, Shi-Xiong Zhang, Guangzhi Li, Yi Luo, and Rongzhi Gu. 2024. Secap: Speech emotion captioning with large language model. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 19323–19331.
- Yoon et al. (2020) Seunghyun Yoon, Subhadeep Dey, Hwanhee Lee, and Kyomin Jung. 2020. Attentive modality hopping mechanism for speech emotion recognition. In ICASSP 2020-2020 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 3362–3366.
- Zellers et al. (2021) Rowan Zellers, Ximing Lu, Jack Hessel, Youngjae Yu, Jae Sung Park, Jize Cao, Ali Farhadi, and Yejin Choi. 2021. Merlot: Multimodal neural script knowledge models. Advances in neural information processing systems 34 (2021), 23634–23651.
- Zhang et al. (2021) Kejun Zhang, Xinda Wu, Xinhang Xie, Xiaoran Zhang, Hui Zhang, Xiaoyu Chen, and Lingyun Sun. 2021. Werewolf-xl: A database for identifying spontaneous affect in large competitive group interactions. IEEE Transactions on Affective Computing 14, 2 (2021), 1201–1214.
- Zhang et al. (2023b) Xiaoqin Zhang, Min Li, Sheng Lin, Hang Xu, and Guobao Xiao. 2023b. Transformer-based multimodal emotional perception for dynamic facial expression recognition in the wild. IEEE Transactions on Circuits and Systems for Video Technology 34, 5 (2023), 3192–3203.
- Zhang et al. (2023a) Yiyuan Zhang, Kaixiong Gong, Kaipeng Zhang, Hongsheng Li, Yu Qiao, Wanli Ouyang, and Xiangyu Yue. 2023a. Meta-transformer: A unified framework for multimodal learning. arXiv preprint arXiv:2307.10802 (2023).
- Zhao et al. ([n. d.]) Bingchen Zhao, Haoqin Tu, Chen Wei, Jieru Mei, and Cihang Xie. [n. d.]. Tuning LayerNorm in Attention: Towards Efficient Multi-Modal LLM Finetuning. In The Twelfth International Conference on Learning Representations.
- Zhao and Patras (2023) Zengqun Zhao and Iliu2022mafwoannis Patras. 2023. Prompting Visual-Language Models for Dynamic Facial Expression Recognition. In BMVC.